OpenStudy (anonymous):

An item is regularly priced at $43 . It is now priced at a discount of 30% off the regular price.

OpenStudy (calculusxy):

Do you wish to find the price after the discount?

OpenStudy (koalagurl1):

$12.90

OpenStudy (koalagurl1):

oops no sorry

OpenStudy (koalagurl1):

$30.10

OpenStudy (koalagurl1):

30% of 43 is 12.9

OpenStudy (welshfella):

-please explain your method @KoalaGurl1

OpenStudy (koalagurl1):

I am @welshfella

OpenStudy (calculusxy):

Assuming that you want to find the price after the discount, what you need to do is: \[$43 - (43 \times .3) = \]

OpenStudy (anonymous):

12.9

OpenStudy (koalagurl1):

so you take away 12.9 from 43

OpenStudy (koalagurl1):

and you get 30.1

OpenStudy (anonymous):

30.1

OpenStudy (koalagurl1):

which is $30.10

OpenStudy (koalagurl1):

Yep :)

OpenStudy (calculusxy):

Or what you can do also is: \($43 \times .7 = x\) I got 0.7 because what it really represents is 70% is what you're paying.